OpenStudy (anonymous):

how do you turn .33333333 into a fraction

OpenStudy (anonymous):

essentially, that would just be 3.33/10

OpenStudy (anonymous):

or if you round it that would be 3/10

OpenStudy (anonymous):

Ah, its the debate of the century, but most math magicians agree, that as a fraction, .33 is 1/3.

OpenStudy (anonymous):

no it comes out to 1/3

OpenStudy (agreene):

1/3 = 0.3333333333333333333333333

OpenStudy (anonymous):

how did u get 1/3

OpenStudy (anonymous):

woops, sorry

OpenStudy (anonymous):

it recurs to infinity so it can be taken as 1/3

OpenStudy (anonymous):

how did u do it?

OpenStudy (anonymous):

?????????????????????????

OpenStudy (agreene):

you can get there from long division

OpenStudy (agreene):

but, most of these types of things people just learn and remember because they come up a lot.

OpenStudy (anonymous):

....

OpenStudy (anonymous):

yea if you divide 3 in to 1 the 3 keeps recurring 3 ) 1.00000000000000 ------------------- 0 . 3333333333333333 .........

OpenStudy (anonymous):

i need to know how to turn the .3333333 into 1/3 not the opposite please

OpenStudy (anonymous):

i've assumed you meant 3 recurring - if the decimal is 0.33333333 - that is it stops in eight pace of decimals then the fraction is 33333333 --------- 100000000

OpenStudy (anonymous):

try this: let x=.333333 then 10x=3.3333 so now 10x = 3.333333... - x = .33333333.... __________________________ 9x = 3 x = 3/9 = 1/3

OpenStudy (lgbasallote):

lol turning decimal to fraction is usually basic grade....algebra???? lol =)))

OpenStudy (lgbasallote):

good method though :p

OpenStudy (anonymous):

or you could also do it by .33333 = 3/10 + 3/100 + 3/1000 ....

OpenStudy (anonymous):

nvm thanx people